OpenStudy (anonymous):

Given f(x)=-2x^3 and g(x)= 4x-5, find g(f(-1))

OpenStudy (anonymous):

PLEASE ANYBODY HELP!

OpenStudy (anonymous):

you should know that big part of math is just plugging it in. plug (-1) into x place. f(-1) = -2(-1)^3=-2(-1)=2 g(f(-1))=g(2)=4(2)-5=8-5=3